### v1

# paper1.bib

Abstract
Quasi-Newton methods form an important class of methods for solving nonlinear optimization problems. In such methods, first order information is used to approximate the second derivative. The aim is to mimic the fast convergence that can be guaranteed by Newton-based methods. In the best case, quasi-Newton methods will far outperform steepest descent and other first order methods, without the computational cost of calculating the exact second derivative. These convergence guarantees hold locally, which follows closely from the fact that if the objective function is strongly convex it can be approximated well by a quadratic function close to the solution. Understanding the performance of quasi-Newton methods on quadratic problems with a symmetric positive definite Hessian is therefore of vital importance. In the classic case, an approximation of the Hessian is updated at every iteration and exact line search is used. It is well known that the algorithm terminates finitely, even when the Hessian approximation is memoryless, i.e. requires only the most recent information. In this paper, we explore the possibilities in which reliance on exact line search and dependence on conjugate search directions can be relaxed, while preserving finite termination properties of quasi-Newton methods on quadratic problems.
Key words. quasi-Newton method, unconstrained quadratic problem, finite termination.
1 Introduction
The main focus of this paper is to study the finite termination properties of quasi-Newton methods on the unconstrained quadratic problem
$\min_{x\in\mathbb{R}^{n}}\frac{1}{2}x^{T}Hx+c^{T}x$ (QP)
where $H=H^{\top}\succ 0$. This is equivalent to solving the linear system of equations $Hx^{*}+c=0$.
Given an initial point $x_{0}$, most methods to solve (QP) in a finite number of iterations generate mutually conjugate search directions $p_{k}$ with respect to $H$. At each iteration, the next iterate is obtained by taking the step
$x_{k+1}=x_{k}+\alpha_{k}p_{k}$
where $\alpha_{k}$ is the optimal step length in the search direction $p_{k}$ found by using exact line search.
In a quasi-Newton method, the search direction $p_{k}$ is found by solving the system $B_{k}p_{k}=-g_{k}$, where $g_{k}=g(x_{k})$ is the gradient at $x_{k}$ and $B_{k}\succ 0$ is some approximation of the Hessian.
The aim of this paper is to investigate the conditions under which a quasi-Newton method can achieve finite termination on (QP) without the need for exact line search. In Section 2 we briefly discuss quasi-Newton and Krylov subspace methods. In Section 3 we introduce the concept of a subspace Newton step. We generalize to a subspace quasi-Newton step in Section 4. Finally we present a quasi-Newton method that can achieve finite termination with the use of arbitrary step sizes in Section 5.
2 Background
In general, to solve (QP) we may find an exact solution in at most $n$ iterations with the use of conjugate search directions and exact line search. Assume that we have fixed some initial point $x_{0}\in\mathbb{R}^{n}$ and $\{p_{0},\dots,p_{n-1}\}$ are a set of conjugate vectors with respect to $H$, i.e. $p_{i}^{T}Hp_{j}=0$ for $i\neq j$. If exact line search is used to calculate the step size
$\alpha_{k}=-\frac{g_{k}^{T}p_{k}}{p_{k}^{T}Hp_{k}},$
[…]
$\min_{x\in\mathbb{R}^{n}}\frac{1}{2}x^{T}Hx+c^{T}x$
subject $\mbox{\rm subject to}\quad x\in x_{0}+\textrm{span }\{p_{0},\dots,p_{k}\}.$
We will focus on Krylov subspace methods, where the search directions span a Krylov subspace as stated in the following definition.
Definition 2.1
Given a vector $b$ and a matrix $A$, the $k$-th Krylov subspace generated by $b$ and $A$ is
$\mathcal{K}_{k}(b,A)=\textrm{span }\{b,Ab,\dots,A^{k-1}b\}.$
Suppose that we fix an initial point $x_{0}$ and search direction $p_{0}=-g_{0}$. Let $\hat{x}_{k}$ denote the minimizer of (QP) in the space $x_{0}+\mathcal{K}_{k}(g_{0},H)$ and $\hat{g}_{k}$ the gradient at that point. The differences $\hat{q}_{k}=\hat{x}_{k+1}-\hat{x}_{k}$ are mutually conjugate with respect to $H$ can be seen as search directions such that exact line search yields the unit step. Additionally $\mathcal{K}_{k}(g_{0},H)$ is spanned by the conjugate search directions $\{\hat{q}_{0},\dots,\hat{q}_{k-1}\}$, or equivalently by the gradients $\{\hat{g}_{0},\dots,\hat{g}_{k-1}\}$.
Once the initial point $x_{0}$ is fixed, a Krylov subspace method with exact line search follows a fixed path until $\hat{x}_{r}=x^{*}$. In the worst case $r=n$, however the algorithm will terminate at iteration $r$, where $r$ is the smallest integer such that $\mathcal{K}_{r}(g_{0},H)=\mathcal{K}_{r+1}(g_{0},H)$.
2.1 Generation of conjugate search directions
In order to solve (QP) using a Krylov subspace method, we must be able to generate search directions that are parallel to $\hat{q}_{k}$. We will denote parallel vectors by $p//p^{\prime}$. When exact line search is used on these parallel directions, the iterates are identical, and as such after $r$ iterations the algorithm will terminate.
One way to generate conjugate directions is to make use of the conjugate Gram-Schmidt process. The method of conjugate gradients (CG) generates these search directions iteratively by setting
$p_{0}=-g_{0}$
$p_{k}=-g_{k}+\frac{g_{k}^{T}Hp_{k-1}}{p_{k-1}^{T}Hp_{k-1}}p_{k-1}.$
For finite termination, it is necessary to use exact line search at each iteration to ensure that the gradients $g_{k}$ are orthogonal to all previous search directions so that the search directions are mutually conjugate.
The Broyden-Fletcher-Goldfarb-Shanno (BFGS) algorithm is a quasi-Newton method, which in the case of a quadratic objective function, will also generate conjugate directions when exact line search is utilized.
Given an initial Hessian approximation $B_{0}\succ 0$, at each iteration a search direction is obtained by solving the system $B_{k}p_{k}=-g_{k}$ and the matrix is updated by
$B_{k+1}=B_{k}-\frac{1}{p_{k}^{T}B_{k}p_{k}}B_{k}p_{k}p_{k}^{T}B_{k}+\frac{1}{p_{k}^{T}Hp_{k}}Hp_{k}p_{k}^{T}H.$
This can be extended to a memoryless variant where the Hessian approximation is built using the identity matrix instead of the previous approximation $B_{k}$, so that
$B_{k+1}=I-\frac{1}{p_{k}^{T}p_{k}}p_{k}p_{k}^{T}+\frac{1}{p_{k}^{T}Hp_{k}}Hp_{k}p_{k}^{T}H.$
The search directions obtained using CG, BFGS and memoryless BFGS all generate parallel search directions $p_{k}//\hat{q}_{k}$, and are therefore Krylov subspace methods.
2.2 Background results
The suggestion of a quasi-Newton methods was first made by Davidon, see [davidon] and then later further developed by Fletcher and Powell, see [fletcher]. For an overview on quasi-Newton methods see [nocedal, Chapter 4].
It has been shown that under certain conditions, $B_{k}$ will generate conjugate directions that are parallel, and with exact line search achieve finite termination in exactly $r$ iterations, see [forsgren2018exact, forsgren2015connection].
On finite termination of quasi-Newton methods, a result from Kolda, see [kolda], shows that finite termination can occur even when the matrix $B_{k}$ is not updated at every iteration if exact line search is used. A similar result from Powell, see [powell1, powell2], proves finite termination using unit steps as long as the update of $B_{k}$ is skipped in every other iteration.
3 Subspace Newton method
Suppose for some initial point $x_{0}$ we consider the $k$-th Krylov subspace $\mathcal{K}_{k}(g_{0},H)$. We may assume without loss of generality that $x_{0}$ is the origin. Then the Krylov space minimizer $\hat{x}_{k}$ will belong to the subspace $\mathcal{K}_{k}(g_{0},H)=\mathcal{K}_{k}(c,H)$.
For any arbitrary point $x\in\mathcal{K}_{k}(c,H)$ there exists a unique $(k-1)-th$ subspace Newton step $p^{N}_{k-1}(x)$ such that $x+p^{N}_{k-1}(x)=\hat{x}_{k}$.
Let a basis $\{s_{0},\dots,s_{k-1}\}$ of $\mathcal{K}_{k}(c,H)$ be chosen. If $g(x)=Hx+c$ is the gradient at the given point $x$, then the $(k-1)$-th subspace Newton step $p^{N}_{k-1}(x)$ can be expressed as
$p^{N}_{k-1}(x)=\left(\!\begin{array}[]{ccccc}s_{k-1}&s_{k-2}&\cdots&s_{0}\end{array}\!\right)\left(\!\begin{array}[]{c}\beta_{k-1}\\ \beta_{k-2}\\ \vdots\\ \beta_{0}\end{array}\!\right),$
for $\beta_{i}$, $i=0,\dots,k$, given by
$\left(\!\begin{array}[]{c}s_{k-1}^{T}\\ s_{k-2}^{T}\\ \vdots\\ s_{0}^{T}\end{array}\!\right)H\left(\!\begin{array}[]{ccccc}s_{k-1}&s_{k-2}&\cdots&s_{0}\end{array}\!\right)\left(\!\begin{array}[]{c}\beta_{k-1}\\ \beta_{k-2}\\ \vdots\\ \beta_{0}\end{array}\!\right)=-\left(\!\begin{array}[]{c}s_{k-1}^{T}\\ s_{k-2}^{T}\\ \vdots\\ s_{0}^{T}\end{array}\!\right)g(x).$
This follows from the fact that $\hat{x}_{k}$ is the unique point that minimizes (QP) over the Krylov subspace, or equivalently that the gradient at the minimizer $\hat{g}_{k}$ is orthogonal to $\mathcal{K}_{k}(c,H)$.
Suppose that instead of an arbitrary basis, we choose a conjugate basis with respect to $H$, $\{q_{0},\dots,q_{k-1}\}$, where $q_{i}//\hat{q}_{i}$. Taking into account that $q_{i}^{T}Hq_{j}=0$ for $i\neq j$, this yields
$\left(\!\begin{array}[]{ccccc}q_{k-1}^{T}Hq_{k-1}&0&0&0\\ 0&q_{k-2}^{T}Hq_{k-2}&0&0\\ 0&0&\ddots&0\\ 0&0&0&q_{0}^{T}Hq_{0}\end{array}\!\right)\left(\!\begin{array}[]{c}\beta_{k-1}\\ \beta_{k-2}\\ \vdots\\ \beta_{0}\end{array}\!\right)=-\left(\!\begin{array}[]{c}q_{k-1}^{T}g(x)\\ q_{k-2}^{T}g(x)\\ \vdots\\ q_{0}^{T}g(x)\end{array}\!\right).$
Therefore it is simple to show that for this basis
$\beta_{i}=-\frac{g(x)^{T}q_{i}}{q_{i}^{T}Hq_{i}}$
for $i=0,\dots,k-1$. We call $\beta_{i}$ the Newton scaling of $q_{i}$ with respect to $x$. Indeed consider $\tilde{x}=x+\sum_{i=0}^{k-1}\beta_{i}q_{i}$. Then for any $j=0,\dots,k-1$
$g(\tilde{x})^{T}q_{j}=\left(g(x)+\sum_{i=0}^{k-1}\beta_{i}Hq_{i}\right)^{T}q_{j}=g(x)^{T}q_{j}+\beta_{j}q_{j}^{T}Hq_{j}=0.$
The $(k-1)$-th subspace Newton step may therefore be recovered using a Newton scaled conjugate basis for $\mathcal{K}_{k}(c,H)$.
Proposition 3.1
Let $x\in\mathcal{K}_{k}(c,H)$ and $p^{N}_{k-1}(x)$ the $k$-th subspace Newton step to $\hat{x}_{k}$. If $q_{k-1}//\hat{q}_{k-1}$ and $\hat{g}_{k}$ is the gradient at $\hat{x}_{k}$, then the $k$-th subspace Newton step $p^{N}_{k}(x)$ is
$p^{N}_{k}(x)$ $=p^{N}_{k-1}(x)+\beta_{k}\hat{g}_{k}+\gamma_{k-1}q_{k-1}$
where
$\beta_{k}=-\frac{\hat{g}_{k}^{T}\hat{g}_{k}q_{k-1}^{T}Hq_{k-1}}{\hat{g}_{k}^{T}H\hat{g}_{k}q_{k-1}^{T}Hq_{k-1}-(\hat{g}_{k}^{T}Hq_{k-1})^{2}}$
$\gamma_{k-1}=\frac{\hat{g}_{k}^{T}\hat{g}_{k}q_{k-1}^{T}H\hat{g}_{k}}{\hat{g}_{k}^{T}H\hat{g}_{k}q_{k-1}^{T}Hq_{k-1}-(\hat{g}_{k}^{T}Hq_{k-1})^{2}}.$
Proof. Since $\hat{x}_{k+1}\in\mathcal{K}_{k+1}(c,H)$, the $k$-th subspace Newton step $p^{N}_{k}(x)$ may be written as a linear combination of $\{\hat{g}_{k},q_{k-1},\dots,q_{0}\}$ where $q_{i}//\hat{q}_{i}$ for $i=0,\dots,k-1$. We assume without loss of generality that $q_{i},i=0,\dots,k-1$ has the Newton scaling with respect to $x$. Taking into account $\hat{g}_{k}^{T}Hq_{i}=0$, $i\leq k-2$, and $q_{i}^{T}\!Hq_{j}=0$, $i\neq j$, we obtain
$\left(\!\begin{array}[]{ccccc}\hat{g}_{k}^{T}H\hat{g}_{k}&\hat{g}_{k}^{T}Hq_{k-1}&0&0&0\\ q_{k-1}^{T}H\hat{g}_{k}&q_{k-1}^{T}Hq_{k-1}&0&0&0\\ 0&0&q_{k-2}^{T}Hq_{k-2}&0&0\\ 0&0&0&\ddots&0\\ 0&0&0&0&q_{0}^{T}Hq_{0}\end{array}\!\right)\left(\!\begin{array}[]{c}\beta_{k}\\ \beta_{k-1}\\ \beta_{k-2}\\ \vdots\\ \beta_{0}\end{array}\!\right)=-\left(\!\begin{array}[]{c}\hat{g}_{k}^{T}g(x)\\ q_{k-1}^{T}g(x)\\ q_{k-2}^{T}g(x)\\ \vdots\\ q_{0}^{T}g(x)\end{array}\!\right).$
Then, $\beta_{i}=1$, $i=0,\dots,k-2$ by the Newton scaling.
Noting that $\hat{g}_{k}=g(x)+Hp^{N}_{k-1}(x)$, we may write $g(x)=\hat{g}_{k}-Hp^{N}_{k-1}(x)$. Since $p^{N}_{k-1}(x)=\sum_{i=0}^{k-1}q_{i}$ we have
$-\hat{g}_{k}^{T}g(x)$ $=$ $-\hat{g}_{k}^{T}(\hat{g}_{k}-Hp^{N}_{k-1}(x))=-\hat{g}_{k}^{T}\hat{g}_{k}+\hat{g}_{k}^{T}Hq_{k-1},$
$-q_{k-1}^{T}g(x)$ $=$ $-q_{k-1}^{T}(\hat{g}_{k}-Hp^{N}_{k-1}(x))=q_{k-1}^{T}Hq_{k-1},$
where the orthogonality of $\hat{g}_{k}$ to $q_{i}$, $i=0,\dots,k-1$ and $Hq_{i}$, $i=0,\dots,k-2$, as well as the conjugacy of the $q_{i}$s have been used.
Then, $\beta_{k}$ and $\beta_{k-1}$ are given by
$\left(\!\begin{array}[]{ccccc}\hat{g}_{k}^{T}H\hat{g}_{k}&\hat{g}_{k}^{T}Hq_{k-1}\\ q_{k-1}^{T}H\hat{g}_{k}&q_{k-1}^{T}Hq_{k-1}\end{array}\!\right)\left(\!\begin{array}[]{c}\beta_{k}\\ \beta_{k-1}\end{array}\!\right)=\left(\!\begin{array}[]{c}-\hat{g}_{k}^{T}\hat{g}_{k}+\hat{g}_{k}^{T}Hq_{k-1}\\ q_{k-1}^{T}Hq_{k-1}\end{array}\!\right).$
Let $\beta_{k-1}=1+\gamma_{k-1}$, then this is reduced to
$\left(\!\begin{array}[]{ccccc}\hat{g}_{k}^{T}H\hat{g}_{k}&\hat{g}_{k}^{T}Hq_{k-1}\\ q_{k-1}^{T}H\hat{g}_{k}&q_{k-1}^{T}Hq_{k-1}\end{array}\!\right)\left(\!\begin{array}[]{c}\beta_{k}\\ \gamma_{k-1}\end{array}\!\right)=\left(\!\begin{array}[]{c}-\hat{g}_{k}^{T}\hat{g}_{k}\\ 0\end{array}\!\right),$
so that
$\left(\!\begin{array}[]{c}\beta_{k}\\ \gamma_{k-1}\end{array}\!\right)$ $=$ $-\frac{\hat{g}_{k}^{T}\hat{g}_{k}}{\hat{g}_{k}^{T}H\hat{g}_{k}q_{k-1}^{T}Hq_{k-1}-(\hat{g}_{k}^{T}Hq_{k-1})^{2}}\left(\!\begin{array}[]{ccccc}q_{k-1}^{T}Hq_{k-1}&-\hat{g}_{k}^{T}Hq_{k-1}\\ -q_{k-1}^{T}H\hat{g}_{k}&\hat{g}_{k}^{T}H\hat{g}_{k}\end{array}\!\right)\left(\!\begin{array}[]{c}1\\ 0\end{array}\!\right)$
$=$ $-\frac{\hat{g}_{k}^{T}\hat{g}_{k}}{\hat{g}_{k}^{T}H\hat{g}_{k}q_{k-1}^{T}Hq_{k-1}-(\hat{g}_{k}^{T}Hq_{k-1})^{2}}\left(\!\begin{array}[]{ccccc}q_{k-1}^{T}Hq_{k-1}\\ -q_{k-1}^{T}H\hat{g}_{k}\end{array}\!\right).$
The $k$-th subspace Newton step to $\hat{x}_{k+1}$ is therefore formed by
$p^{N}_{k}(x)=\sum_{i=0}^{k-1}q_{i}+\beta_{k}\hat{g}_{k}+\gamma_{k-1}q_{k-1}=p^{N}_{k-1}(x)+\beta_{k}\hat{g}_{k}+\gamma_{k-1}q_{k-1}$
where $\beta_{k}\hat{g}_{k}+\gamma_{k-1}q_{k-1}=\hat{q}_{k}$.
Note that $\beta_{k}$ and $\gamma_{k-1}q_{k-1}$ are independent of the scaling of $q_{k-1}$, thus the formula for the $k$-th subspace Newton step is valid for any $q_{k-1}//\hat{q}_{k-1}$.
If we consider the case when $x=\hat{x}_{k}$, then $p^{N}_{k-1}(\hat{x}_{k})=0$. Therefore
$p^{N}_{k}(\hat{x}_{k})=-\beta_{k}\left(-\hat{g}_{k}-\frac{\hat{g}_{k}^{T}Hq_{k-1}}{q_{k-1}^{T}Hq_{k-1}}q_{k-1}\right),$
so we recover the method of conjugate gradients equipped with exact line search.
4 Quasi-Newton subspace approximation
To recover the $k$-th subspace Newton step from $x\in\mathcal{K}_{k}(c,H)$ to $\hat{x}_{k+1}$ as presented in the previous section, we require knowledge on how $H$ acts on $K_{k+1}(c,H)$. In particular we must have sufficient information to calculate the Newton scaling of the conjugate vectors $q_{0},\dots,q_{k-1}$ with respect to $x$ as well as the scalars $\beta_{k},\gamma_{k-1}$.
Suppose instead we simply have a quasi-Newton approximation $B_{k}\succ 0$ such that $B_{k}q_{i}=Hq_{i}$ for $i=0,\dots,k-1$. Then the only term that is unknown is $\hat{g}_{k}^{T}H\hat{g}_{k}$, which must be replaced by the approximation $\hat{g}_{k}^{T}B_{k}\hat{g}_{k}$.
If such an approximation is used, then we may construct a step $\tilde{p}_{k}(x)=p_{k-1}^{N}(x)+\delta_{k}\hat{q}_{k}$ by solving the system
$\left(\!\begin{array}[]{ccccc}\hat{g}_{k}^{T}B_{k}\hat{g}_{k}&\hat{g}_{k}^{T}Hq_{k-1}\\ q_{k-1}^{T}H\hat{g}_{k}&q_{k-1}^{T}Hq_{k-1}\end{array}\!\right)\left(\!\begin{array}[]{c}\tilde{\beta}_{k}\\ \tilde{\gamma}_{k-1}\end{array}\!\right)=\left(\!\begin{array}[]{c}-\hat{g}_{k}^{T}\hat{g}_{k}\\ 0\end{array}\!\right),$
so that
$\left(\!\begin{array}[]{c}\tilde{\beta}_{k}\\ \tilde{\gamma}_{k-1}\end{array}\!\right)=-\frac{\hat{g}_{k}^{T}\hat{g}_{k}}{\hat{g}_{k}^{T}B_{k}\hat{g}_{k}q_{k-1}^{T}Hq_{k-1}-(\hat{g}_{k}^{T}Hq_{k-1})^{2}}\left(\!\begin{array}[]{ccccc}q_{k-1}^{T}Hq_{k-1}\\ -q_{k-1}^{T}H\hat{g}_{k}\end{array}\!\right)=\delta_{k}\left(\!\begin{array}[]{c}\beta_{k}\\ \gamma_{k-1}\end{array}\!\right),$
for
$\delta_{k}=\frac{\hat{g}_{k}^{T}H\hat{g}_{k}q_{k-1}^{T}Hq_{k-1}-(\hat{g}_{k}^{T}Hq_{k-1})^{2}}{\hat{g}_{k}^{T}B_{k}\hat{g}_{k}q_{k-1}^{T}Hq_{k-1}-(\hat{g}_{k}^{T}Hq_{k-1})^{2}}.$
However, to do this we require explicit knowledge of the gradient $\hat{g}_{k}$. To avoid this, we will build an approximation $B_{k}$ which acts on the Krylov subspace $\mathcal{K}_{k}(c,H)$ in such a way that for any $x\in\mathcal{K}_{k}(c,H)$, solving the system of equations $B_{k}p_{k}(x)=-g(x)$ yields a step $p_{k}(x)=p_{k-1}^{N}(x)+\delta_{k}\hat{q}_{k}$ without requiring knowledge of $\hat{g}_{k}$
Assume that $B_{k}$ is given by
$B_{k}=\sigma_{k}(I-Q_{k-1}(Q_{k-1}^{T}Q_{k-1})^{-1}Q_{k-1}^{T})+HQ_{k-1}(Q_{k-1}^{T}HQ_{k-1})^{-1}Q_{k-1}^{T}H,$ (4.1)
where $Q_{k-1}=(q_{k-1}\ q_{k-2}\ \dots\ q_{0})$ for $q_{i}//\hat{q}_{i},i=0,\dots,k-1$ and $\sigma_{k}>0$. Then $B_{k}$ acts on $\mathcal{K}_{k}(c,H)$ by design.
Proposition 4.1
Let $x\in\mathcal{K}_{k}(c,H)$ and $B_{k}$ be given by (4.1). Then $B_{k}p_{k}=-g(x)$ has a unique solution given by
$p_{k}(x)$ $=\frac{1}{\sigma_{k}}\left(-\hat{g}_{k}+\frac{\hat{g}_{k}^{T}\!Hq_{k-1}}{q_{k-1}^{T}Hq_{k-1}}q_{k-1}\right)+\sum_{i=0}^{k-1}\left(-\frac{q_{i}^{T}\!g(x)}{q_{i}^{T}\!Hq_{i}}\right)q_{i}.$ (4.2)
Additionally there exists some unique $\sigma_{k}$ such that $p_{k}(x)=p^{N}_{k}(x)$.
Proof. Since $\hat{g}_{k}$ is the gradient at $\hat{x}_{k}$, it must belong to the null space of $Q_{k-1}$. Additionally, by construction the term
$q_{k}=-\hat{g}_{k}+\frac{\hat{g}_{k}^{T}\!Hq_{k-1}}{q_{k-1}^{T}Hq_{k-1}}q_{k-1}$
is conjugate to the vectors $q_{0},\dots,q_{k-1}$ with respect to $H$ and belongs to $K_{k+1}(c,H)$. Therefore $B_{k}q_{k}=\sigma_{k}\hat{g}_{k}$.
Additionally, given that
$p_{k-1}^{N}(x)=\sum_{i=0}^{k-1}\left(-\frac{q_{i}^{T}\!g(x)}{q_{i}^{T}\!Hq_{i}}\right)q_{i}$
we may write $g(x)=\hat{g}_{k}-Hp_{k-1}^{N}(x)$.
Applying $B_{k}$ to (4.2) therefore yields the desired result
$B_{k}p_{k}(x)$ $=-\hat{g}_{k}+\sum_{i=0}^{k-1}\left(-\frac{q_{i}^{T}\!g(x)}{q_{i}^{T}\!Hq_{i}}\right)Hq_{i}$
$=-g(x).$
For $p_{k}(x)=p_{k}^{N}(x)$ we require that each $q_{i},i=0,\dots k$ has the Newton scaling with respect to $x$, or equivalently that $q_{k}=\sigma_{k}\hat{q}_{k}$. In other words
$\sigma_{k}=-\frac{q_{k}^{T}\!Hq_{k}}{q_{k}^{T}\!g(x)}$
is the unique value for which $B_{k}$ recovers the $k$-th subspace Newton step.
The approximation $B_{k}$ given by (4.1) is independent of the choice of $x\in\mathcal{K}_{k}(c,H)$. As $B_{k}$ acts on each $q_{0},\dots,q_{k-1}$ as $H$, solving the system $B_{k}p_{k}(x)=-g(x)$ can recover the $(k-1)$-th subspace Newton step by appropriately rescaling the conjugate vectors.
Suppose that instead of using a full basis for $\mathcal{K}_{k}(c,H)$ we limit ourselves to $P_{k-1}=(p^{N}_{k-1}(x)\ q_{k-1})$. Unlike $Q_{k-1}$ which is made of a full basis of conjugate vectors, $P_{k-1}$ requires the $(k-1)$-th subspace Newton step from $x$ to $\hat{x}_{k}$, and is therefore dependent on $x$. We may then build an approximation $B_{k}(x)$ using $P_{k-1}$ by
$B_{k}(x)=\sigma_{k}(I-P_{k-1}(P_{k-1}^{T}P_{k-1})^{-1}P_{k-1}^{T})+HP_{k-1}(P_{k-1}^{T}HP_{k-1})^{-1}P_{k-1}^{T}H$ (4.3)
which will act on a two dimensional subspace of $\mathcal{K}_{k}(c,H)$ as $H$. In particular $B_{k}(x)$ holds information on how $H$ acts on the subspace Newton step from $x$ to $\hat{x}_{k}$ and the vector $q_{k}//\hat{q}_{k}$, which is sufficient to recover the step $p_{k}(x)$.
Proposition 4.2
Let $x\in\mathcal{K}_{k}(c,H)$ and $B_{k}(x)$ be given by (4.3). Then $B_{k}(x)p_{k}(x)=-g(x)$ has a unique solution given by
$p_{k}(x)$ $=\frac{1}{\sigma_{k}}\left(-\hat{g}_{k}+\frac{\hat{g}_{k}^{T}\!Hq_{k-1}}{q_{k-1}^{T}Hq_{k-1}}q_{k-1}\right)+p_{k-1}^{N}(x).$ (4.4)
which coincides with the solution calculated by the approximation $B_{k}$ given by (4.1).
Proof. By the construction given in (4.3), $B_{k}(x)$ will act on $p_{k-1}^{N}(x)$ and $q_{k-1}$ as $H$. Additionally
$q_{k}=-\hat{g}_{k}+\frac{\hat{g}_{k}^{T}\!Hq_{k-1}}{q_{k-1}^{T}Hq_{k-1}}q_{k-1}$
must be conjugate to $q_{i}$ with respect to $H$ for $i=0,\dots,k-1$ and $q_{k}//\hat{q}_{k}$.
Therefore, just as $B_{k}$ given by (4.1) acts on $q_{k}$, since $q_{k}$ in the null space of $HP_{k-1}$ and $\hat{g}_{k}$ is in the null space of $P_{k-1}$, we have that $B_{k}(x)q_{k}=\sigma_{k}\hat{g}_{k}$. Applying $B_{k}(x)$ to (4.4) yields
$B_{k}p_{k}(x)$ $=-\hat{g}_{k}+Hp^{N}_{k-1}(x)$
$=-g(x).$
Given that
$p_{k-1}^{N}(x)=\sum_{i=0}^{k-1}\left(-\frac{q_{i}^{T}\!g(x)}{q_{i}^{T}\!Hq_{i}}\right)q_{i}$
we have that the equations (4.2) and (4.4) coincide.
5 An exact Quasi-Newton method without exact line search
If we are given an arbitrary point $x\in\mathcal{K}_{k}(c,H)$ then generating the $k$-th subspace Newton step or quasi-Newton step requires knowledge of the vectors $q_{i}//\hat{q}_{i}$ for $i=0,\dots k-1$. In addition it is also vital that we have information on how $H$ acts on these vectors to recover the Newton scaling. This is in general not information that is readily available.
We propose instead an iterative learning algorithm which generates subspace quasi-Newton steps using arbitrary step sizes. After a finite number of iterations, the algorithm will generate the Newton step, which will lead to termination if a unit step is used.
The step size $\alpha_{k}$ can be chosen freely. Theoretically, if we have explicit knowledge of $H$ then even a zero step will lead to generation of the Newton step as iteratively new conjugate directions are calculated and then Newton scaled with respect to the fixed point.
Without explicit knowledge of $H$ it is vital to take a nonzero step in order to evaluate a new gradient. We may then use the difference of gradients to learn how $H$ acts on the step taken. This method is described in Algorithm 1.
Algorithm 1 Quasi-Newton subspace method
$x_{0}\in\mathbb{R}^{n},B_{0}=I,p^{N}_{-1}=0$
while $k=0,1,\dots$ do
     $g_{k}\leftarrow Hx_{k}+c$
     Solve $B_{k}p_{k}=-g_{k}$
     Choose step size $\alpha_{k}\neq 0$
     $x_{k+1}\leftarrow x_{k}+\alpha_{k}p_{k}$
     $q_{k}\leftarrow p_{k}-p^{N}_{k-1}$
     $p^{N}_{k}\leftarrow(1-\alpha_{k})p^{N}_{k-1}-\left(\frac{g_{k}^{T}q_{k}}{q_{k}^{T}Hq_{k}}+\alpha_{k}\right)q_{k}$
     if $q_{k}//p^{N}_{k}$ then
         $P=\begin{bmatrix}q_{k}\end{bmatrix}$
     else
         $P=\begin{bmatrix}p^{N}_{k}&q_{k}\end{bmatrix}$
     end if
     Choose some $\sigma_{k}>0$
     $B_{k+1}=\sigma_{k}(I-P(P^{T}P)^{-1}P^{T})+HP(P^{T}HP)^{-1}P^{T}H$
     $k\leftarrow k+1$
end while
Theorem 5.1
For a fixed initial point $x_{0}$, consider iteration $k$ of Algorithm 1, such that $\alpha_{i}\neq 0,\sigma_{i}>0$ for $i=0,\dots,k-1$. Let $B_{k}$ be the matrix generated by the algorithm. Then $B_{k}=B(x_{k})$ as given by (4.3). Additionally, the algorithm will terminate if and only if $\alpha_{k}=1$ for some $k\geq r$, where $r$ is the dimension of the largest Krylov subspace generated by $g_{0}$ and $H$.
Proof.
To show that the Hessian approximation $B_{k}$ generated by Algorithm 1 is exactly $B_{k}(x_{k})$ given by (4.3), it is sufficient to show that $q_{k}//\hat{q}_{k}$ and that $p_{k}^{N}=p_{k}^{N}(x_{k+1})$. We proceed by induction. For $k=0$, clearly $q_{0}=-g(x_{0})$ and $p_{0}^{N}$ is Newton scaled with respect to $x_{1}$.
Assume that $q_{k-1}//\hat{q}_{k-1}$ and $p_{k-1}^{N}=p_{k-1}^{N}(x_{k})$. Then the matrix $B_{k}=B_{k}(x_{k})$ so the search direction $p_{k}$ is exactly that given by Proposition 4.2. Therefore $q_{k}=p_{k}-p_{k-1}^{N}(x_{k})//\hat{q}_{k}$.
Additionally consider the gradient $g(x_{k+1}+p_{k}^{N})$. For $q_{i},i\leq k-1$
$g(x_{k+1}+p_{k}^{N})^{T}q_{i}$ $=(g_{k}+\alpha_{k}Hp_{k-1}^{N}(x_{k})+(1-\alpha_{k})Hp_{k-1}^{N}(x_{k}))^{T}q_{i}$
$=(g_{k}+Hp_{k-1}^{N}(x_{k}))^{T}q_{i}$
$=0.$
For $q_{k}$, we may use the fact that $q_{k}$ is conjugate to $p_{k-1}^{N}(x_{k})$ with respect to $H$ to show that
$g(x_{k+1}+p_{k}^{N})^{T}q_{k}$ $=(g_{k}+\alpha_{k}Hq_{k}-(\frac{g_{k}^{T}q_{k}}{q_{k}^{T}Hq_{k}}+\alpha_{k})Hq_{k})^{T}q_{k}$
$=(g_{k}-\frac{g_{k}^{T}q_{k}}{q_{k}^{T}Hq_{k}}Hq_{k})^{T}q_{k}$
$=0.$
For $i\geq r$, the search direction $p_{i}=p_{i}^{N}$ generated by Algorithm 1 is therefore the Newton step from $x_{i+1}$ to $x^{*}$. Termination is therefore possible if and only if $\alpha_{i}=1$ for some $i>r$.
In the case that exact line search is used at each iteration of Algorithm 1, we recover BFGS with $B_{0}=\sigma_{k}I$ at iteration $k$. In such a case it is well known that $B_{k+1}$ built by the most recent conjugate vector $q_{k}$ is sufficient to guarantee finite termination.
In the case that $\alpha_{k}=1$ always, it is also sufficient to build $B_{k+1}$ using only $q_{k}$. Additionally, when unit step is always used, the algorithm will always terminate in exactly $r$ or $r+1$ iterations, as stated in the corollary below.
Corollary 5.1
Assume that the step $\alpha_{k}=1$ is chosen at each iteration of Algorithm 1. Then it is sufficient to build $B_{k}$ using $q_{k}$. There exists a unique value of $\sigma_{r-1}$ such that the algorithm terminates in exactly $r$ iterations. For all other values of $\sigma_{r-1}$ the algorithm will terminate in exactly $r+1$ iterations.
Proof. If $\alpha_{k}=1$ for each iteration of Algorithm 1, then by definition
$p^{N}_{k}$ $=(1-\alpha_{k})p^{N}_{k-1}-\left(\frac{g_{k}^{T}q_{k}}{q_{k}^{T}Hq_{k}}+\alpha_{k}\right)q_{k}$
$=-\left(\frac{g_{k}^{T}q_{k}}{q_{k}^{T}Hq_{k}}+\alpha_{k}\right)q_{k}//q_{k}.$
Therefore one vector is sufficient to build $B_{k}$.
As shown in Proposition 4.2, there exists a unique scalar $\sigma_{r-1}$ such that $B_{r}$ will generate the Newton step. In this case the algorithm will terminate if unit step is used. Otherwise, $B_{r+1}$ acts as $H$ on the gradient $g_{r+1}$ and therefore the search direction generated will be the Newton step.
Hence, using the correct vectors to create the Hessian approximation $B_{k}$ can lead to termination on (QP) by calculating the Newton step in a finite number of iterations.
6 Conclusion
In this paper we show sufficient conditions in which to construct a Hessian approximation $B_{k}$ such that exact line search is not necessary to find the Newton step in a finite number of iterations.
The main contribution of this paper is a learning algorithm which can gather information for the Hessian approximation regardless of step sizes that are used as shown in Theorem 5.1. In addition, we show in Corollary 5.1, that without line search, unit step sizes will lead to termination in exactly $r$ or $r+1$ iterations.
Our focus has been to explore what are the requirements for a quasi-Newton method to be exact beyond the usual expectations of conjugate search directions and exact line search. We hope that this can lead to a deeper understanding on the behavior of quasi-Newton methods on general nonlinear problems.
[…]