### v2

# paper1.bib

Abstract
Quasi-Newton methods form an important class of methods for solving nonlinear optimization problems. In such methods, first order information is used to approximate the second derivative. The aim is to mimic the fast convergence that can be guaranteed by Newton-based methods. In the best case, quasi-Newton methods will far outperform steepest descent and other first order methods, without the computational cost of calculating the exact second derivative. These convergence guarantees hold locally, which follows closely from the fact that, if the objective function is strongly convex, it can be approximated well by a quadratic function close to the solution. Understanding the performance of quasi-Newton methods on quadratic problems with a symmetric positive definite Hessian is therefore of vital importance. In the classic case, an approximation of the Hessian is updated at every iteration and exact line search is used. It is well known that the algorithm terminates finitely, even when the Hessian approximation is memoryless, i.e. requires only the most recent information. In this paper, we explore the possibilities in which reliance on exact line search and dependence on conjugate search directions can be relaxed, while preserving finite termination properties of quasi-Newton methods on quadratic problems. We show that it suffices to create a memoryless quasi-Newton matrix based on two vectors to give ability to compute a Newton direction within a finite number of iterations, independent of step lengths. It is unnecessary for the quasi-Newton approximation to act as the Hessian on the full space.
Key words. quasi-Newton method, unconstrained quadratic problem, finite termination.
1 Introduction
The main focus of this paper is to study the finite termination properties of quasi-Newton methods on the unconstrained quadratic problem
$\min_{x\in\mathbb{R}^{n}}\frac{1}{2}x^{T}Hx+c^{T}x$ (QP)
where $H=H^{\top}\succ 0$. This is equivalent to solving the linear system of equations $Hx^{*}+c=0$.
Given an initial point $x_{0}$, (QP) may be solved in a finite number of iterations by an iterative method that generates mutually conjugate search directions $p_{k}$ with respect to $H$. At each iteration, the next iterate is obtained by taking the step
$x_{k+1}=x_{k}+\alpha_{k}p_{k}$
where $\alpha_{k}$ is the optimal step length in the search direction $p_{k}$ found by using exact line search.
In a quasi-Newton method, the search direction $p_{k}$ is found by solving the system $B_{k}p_{k}=-g_{k}$, where $g_{k}=g(x_{k})$ is the gradient at $x_{k}$ and $B_{k}=B_{k}^{T}\succ 0$ is some approximation of the Hessian $H$.
The aim of this paper is to investigate the conditions under which a quasi-Newton method can achieve finite termination on (QP) without the need for exact line search. In Section 2 we briefly discuss Krylov subspaces and Krylov subspace methods such as the method of conjugate gradients as well as quasi-Newton methods. In Section 3 we introduce the concept of a subspace Newton step within a given Krylov subspace as well as an expanding Krylov subspace. We generalize to a subspace quasi-Newton step in Section 4, showing that it suffices to create a quasi-Newton matrix based on at most two vectors. Finally we present an iterative memoryless quasi-Newton method based on at most two vectors, in addition to the gradient, that will compute the Newton step in a finite number of iterations with the use of arbitrary step sizes in Section 5.
2 Background
In general, to solve (QP) we may find an exact solution in at most $n$ iterations with the use of conjugate search directions and exact line search. Given a point $x\in\mathbb{R}^{n}$ and a search direction $p$, exact line search can be used to calculate the optimal step length
$\alpha=-\frac{g(x)^{T}p}{p^{T}Hp}.$
[…]
$\min_{x\in\mathbb{R}^{n}}\frac{1}{2}x^{T}Hx+c^{T}x$
subject $\mbox{\rm subject to}\quad x\in x_{0}+\textrm{span }\{p_{0},\dots,p_{k}\}.$
The focus of this paper will be on Krylov subspace methods, where the search directions span a Krylov subspace as stated in the following definition.
Definition 2.1
Given a vector $b$ and a matrix $A$, the $k$-th Krylov subspace generated by $b$ and $A$ is
$\mathcal{K}_{k}(b,A)=\textrm{span }\{b,Ab,\dots,A^{k-1}b\}.$
Suppose that we fix an initial point $x_{0}$, search direction $p_{0}=-g_{0}$ and consider the Krylov subspaces $\mathcal{K}_{k}(g_{0},H)$. Any Krylov subspace method will follow a fixed path to minimizers over expanding affine spaces until termination occurs in a finite number of iterations.
Definition 2.2
Given $x_{0}\in\mathbb{R}^{n}$, let $r$ be the smallest integer such that $\mathcal{K}_{r}(g_{0},H)=\mathcal{K}_{r+1}(g_{0},H)$. For $k\leq r$, denote the minimizer of (QP) over the affine space $x_{0}+\mathcal{K}_{k}(g_{0},H)$ as $\hat{x}_{k}$. The gradient at the minimizer is denoted by $\hat{g}_{k}$ and the search direction generated by any Krylov subspace method is parallel to the difference $\hat{q}_{k-1}=\hat{x}_{k}-\hat{x}_{k-1}$.
The search directions $\hat{q}_{0},\dots,\hat{q}_{k-1}$ are mutually conjugate with respect to $H$ and form a basis for $\mathcal{K}_{k}(g_{0},H)$. Additionally, the gradients $\hat{g}_{0},\dots,\hat{g}_{k-1}$ form an orthogonal basis for the same subspace. For $r$ as defined above, the minimizer $\hat{x}_{r}$ is exactly the global minimizer $x^{*}$ of (QP). For a more comprehensive overview of Krylov subspace methods, see, e.g [nocedal, Chapter 5], [Saa95, Chapter 6].
2.1 Generation of conjugate search directions using exact line search
In order to solve (QP) using a Krylov subspace method, search directions that are parallel to $\hat{q}_{k}$ must be generated. Parallel vectors will be denoted by $p//p^{\prime}$. When exact line search is used on these parallel directions, the iterates are identical, and as such after $r$ iterations the algorithm will terminate.
One way to generate conjugate directions is to make use of the conjugate Gram-Schmidt process. The method of conjugate gradients (CG) generates these search directions iteratively by setting
$p_{0}=-g_{0},\quad p_{k}=-g_{k}+\frac{g_{k}^{T}Hp_{k-1}}{p_{k-1}^{T}Hp_{k-1}}p_{k-1}=-g_{k}+\frac{g_{k}^{T}g_{k}}{g_{k-1}^{T}g_{k-1}}p_{k-1}.$
For finite termination, it is necessary to use exact line search at each iteration to ensure that the gradients $g_{k}$ are orthogonal to all previous search directions so that the search directions are mutually conjugate. For a detailed description of the methods of conjugate gradients, see, e.g. Saad [Saa95, Chapter 6.7].
2.2 Quasi-Newton methods
The Broyden-Fletcher-Goldfarb-Shanno (BFGS) algorithm is a quasi-Newton method, which in the case of a quadratic objective function, will also generate conjugate directions when exact line search is utilized. Given an initial Hessian approximation $B_{0}\succ 0$, at each iteration a search direction is obtained by solving the system $B_{k}p_{k}=-g_{k}$ and the matrix is updated by
$B_{k+1}=B_{k}-\frac{1}{p_{k}^{T}B_{k}p_{k}}B_{k}p_{k}p_{k}^{T}B_{k}+\frac{1}{p_{k}^{T}Hp_{k}}Hp_{k}p_{k}^{T}H.$
This can be extended to a memoryless variant where the Hessian approximation is built using the identity matrix instead of the previous approximation $B_{k}$, so that
$B_{k+1}=I-\frac{1}{p_{k}^{T}p_{k}}p_{k}p_{k}^{T}+\frac{1}{p_{k}^{T}Hp_{k}}Hp_{k}p_{k}^{T}H.$
The different methods CG, BFGS and memoryless BFGS all generate parallel search directions $p_{k}//\hat{q}_{k}$ if exact line search is used, and are therefore Krylov subspace methods [huang].
The suggestion of a quasi-Newton methods was first made by Davidon [davidon] and then later further developed by Fletcher and Powell [fletcher]. For an overview on quasi-Newton methods, see, e.g., [nocedal, Chapter 4], [GMW19, Chapter 4.5] and [Fle81, Chapter 3.5]. There is a vast literature on quasi-Newton methods, which is beyond the scope of the current manuscript. We will point to the BFGS method and the memoryless BFGS method, which are of relevance for our results.
It has been shown that under certain conditions, a quasi-Newton matrix $B_{k}$ will generate conjugate directions that are parallel, and with exact line search achieve finite termination in exactly $r$ iterations [forsgren2018exact, forsgren2015connection].
On finite termination of quasi-Newton methods, a result from Kolda, O’Leary and Nazareth [kolda], shows that finite termination can occur even when the quasi-Newton matrix $B_{k}$ is not updated at every iteration if exact line search is used. A similar result from Powell [powell1, powell2], proves finite termination using unit steps as long as the update of the quasi-Newton matrix $B_{k}$ is skipped in every other iteration.
3 Subspace Newton method
Suppose for some initial point $x_{0}$ we consider the Krylov subspace $\mathcal{K}_{k}(g_{0},H)$. Then the Krylov space minimizer $\hat{x}_{k}$ will belong to the affine space $x_{0}+\mathcal{K}_{k}(g_{0},H)$. For any arbitrary point $x\in x_{0}+\mathcal{K}_{k}(g_{0},H)$ there exists a unique subspace Newton step $p^{N}_{k-1}(x)\in\mathcal{K}_{k}(g_{0},H)$ such that $x+p^{N}_{k-1}(x)=\hat{x}_{k}$. Suppose $S=\{s_{0},\dots,s_{k-1}\}$ has columns that form a basis of $\mathcal{K}_{k}(g_{0},H)$. If $g(x)=Hx+c$ is the gradient at the given point $x$, then the subspace Newton step $p^{N}_{k-1}(x)$ can be expressed as
$p^{N}_{k-1}(x)=S\beta,$
for $\beta$ given by
$S^{T}HS\beta=-S^{T}g(x).$
This follows from the fact that $\hat{x}_{k}$ is the unique point that minimizes (QP) over the affine space $x_{0}+\mathcal{K}_{k}(g_{0},H)$, or equivalently that the gradient at the minimizer $\hat{g}_{k}$ is orthogonal to $\mathcal{K}_{k}(g_{0},H)$. The subspace Newton step will be fundamental in our analysis in that the relative change of the directional derivatives of the gradient along the subspace Newton step with respect to all conjugate directions can be characterized by one steplength $\alpha$, as stated in the following lemma.
Lemma 3.1
For $k\leq r$, let $q_{0},\dots,q_{k-1}$ be a basis for $\mathcal{K}_{k}(g_{0},H)$ where $q_{i}//\hat{q}_{i}$. For any $x\in x_{0}+\mathcal{K}_{k}(g_{0},H)$, the subspace Newton step $p^{N}_{k-1}(x)$ to $\hat{x}_{k}$ is
$p_{k-1}^{N}(x)=\sum_{i=0}^{k-1}-\frac{g(x)^{T}q_{i}}{q_{i}^{T}Hq_{i}}q_{i}.$
Additionally, for any $\alpha\in\mathbb{R}$
$g(x+\alpha p_{k-1}^{N}(x))^{T}q_{j}=(1-\alpha)g(x)^{T}q_{j},\quad j=0,\dots,k-1.$
Proof. Since $q_{i}//\hat{q}_{i},i=0,\dots,k-1$, the basis $q_{0},\dots,q_{k-1}$ is mutually conjugate with respect to $H$. We may therefore represent the subspace Newton step as $p^{N}_{k-1}(x)=\sum_{i=0}^{k-1}\beta_{i}q_{i}$ where $\beta_{i},i=0,\dots,k-1$ are the solutions to the system
$\left(\!\begin{array}[]{ccccc}q_{k-1}^{T}Hq_{k-1}&0&0&0\\ 0&q_{k-2}^{T}Hq_{k-2}&0&0\\ 0&0&\ddots&0\\ 0&0&0&q_{0}^{T}Hq_{0}\end{array}\!\right)\left(\!\begin{array}[]{c}\beta_{k-1}\\ \beta_{k-2}\\ \vdots\\ \beta_{0}\end{array}\!\right)=-\left(\!\begin{array}[]{c}q_{k-1}^{T}g(x)\\ q_{k-2}^{T}g(x)\\ \vdots\\ q_{0}^{T}g(x)\end{array}\!\right).$
Therefore it is straightforward to show that
$\beta_{i}=-\frac{g(x)^{T}q_{i}}{q_{i}^{T}Hq_{i}},\qquad i=0,\dots,k-1$
and the subspace Newton step is
$p_{k-1}^{N}(x)=\sum_{i=0}^{k-1}-\frac{g(x)^{T}q_{i}}{q_{i}^{T}Hq_{i}}q_{i}.$
Additionally, for any $\alpha\in\mathbb{R}$ and $j=0,\dots,k-1$,
$g(x+\alpha p_{k-1}^{N}(x))^{T}\!q_{j}$ $=(g(x)-\alpha\sum_{i=0}^{k-1}\frac{g(x)^{T}\!q_{i}}{q_{i}^{T}Hq_{i}}Hq_{i})^{T}q_{j}$
$=g(x)^{T}\!q_{j}-\alpha\frac{g(x)^{T}\!q_{j}}{q_{j}^{T}Hq_{j}}q_{j}^{T}Hq_{j}=(1-\alpha)g(x)^{T}q_{j},$
where the conjugacy of the vectors $q_{0},\dots,q_{k-1}$ has been used.
For a point $x\in x_{0}+\mathcal{K}_{k}(g_{0},H)$, the subspace Newton step to $\hat{x}_{k}$ may therefore be recovered using a Newton scaled conjugate basis for $\mathcal{K}_{k}(g_{0},H)$. Once the step $p^{N}_{k-1}(x)$ is known explicitly, it is possible to recover $\hat{x}_{k}$ and therefore the gradient $\hat{g}_{k}$. Since $\hat{g}_{k}$ is orthogonal to $\mathcal{K}_{k}(g_{0},H)$, the vectors $\{\hat{g}_{k},q_{k-1},\dots,q_{0}\}$ form a basis for $\mathcal{K}_{k+1}(g_{0},H)$. Using such a basis, we may construct the subspace Newton step $p^{N}_{k}(x)$ to the minimizer $\hat{x}_{k+1}\in x_{0}+\mathcal{K}_{k+1}(g_{0},H)$ for any $x\in x_{0}+\mathcal{K}_{k}(g_{0},H)$.
Proposition 3.1
For $k<r$, let $x\in x_{0}+\mathcal{K}_{k}(g_{0},H)$ and let $p^{N}_{k-1}(x)$ be the subspace Newton step to $\hat{x}_{k}$. If $q_{k-1}//\hat{q}_{k-1}$ and $\hat{g}_{k}$ is the gradient at $\hat{x}_{k}$, then the subspace Newton step $p^{N}_{k}(x)$ to $\hat{x}_{k+1}$ is
$p^{N}_{k}(x)$ $=\hat{q}_{k}+p_{k-1}^{N}(x)$
with
$\hat{q}_{k}=\frac{1}{\hat{\sigma}_{k}}\left(-\hat{g}_{k}+\frac{\hat{g}_{k}^{T}Hq_{k-1}}{q_{k-1}^{T}Hq_{k-1}}q_{k-1}\right)\text{and }\hat{\sigma}_{k}=\frac{\hat{g}_{k}^{T}H\hat{g}_{k}q_{k-1}^{T}Hq_{k-1}-(\hat{g}_{k}^{T}Hq_{k-1})^{2}}{\hat{g}_{k}^{T}\hat{g}_{k}q_{k-1}^{T}Hq_{k-1}}.$
Proof. Since $\hat{x}_{k+1}\in x_{0}+\mathcal{K}_{k+1}(g_{0},H)$, the subspace Newton step $p^{N}_{k}(x)$ may be written as a linear combination of $\hat{g}_{k},q_{k-1},\dots,q_{0}$ where $q_{i}//\hat{q}_{i}$ for $i=0,\dots,k-1$ which forms a basis for $\mathcal{K}_{k+1}(g_{0},H)$. Taking into account that $\hat{g}_{k}^{T}Hq_{i}=0$, $i\leq k-2$, and $q_{i}^{T}\!Hq_{j}=0$, $i\neq j$, yields
$\left(\!\begin{array}[]{ccccc}\hat{g}_{k}^{T}H\hat{g}_{k}&\hat{g}_{k}^{T}Hq_{k-1}&0&0&0\\ q_{k-1}^{T}H\hat{g}_{k}&q_{k-1}^{T}Hq_{k-1}&0&0&0\\ 0&0&q_{k-2}^{T}Hq_{k-2}&0&0\\ 0&0&0&\ddots&0\\ 0&0&0&0&q_{0}^{T}Hq_{0}\end{array}\!\right)\left(\!\begin{array}[]{c}\beta_{k}\\ \beta_{k-1}\\ \beta_{k-2}\\ \vdots\\ \beta_{0}\end{array}\!\right)=-\left(\!\begin{array}[]{c}\hat{g}_{k}^{T}g(x)\\ q_{k-1}^{T}g(x)\\ q_{k-2}^{T}g(x)\\ \vdots\\ q_{0}^{T}g(x)\end{array}\!\right).$
Then,
$\beta_{i}=-\frac{g(x)^{T}q_{i}}{q_{i}^{T}Hq_{i}},\qquad i=0,\dots,k-2$
is the Newton scaling with respect to $x$.
Noting that $\hat{g}_{k}=g(x)+Hp^{N}_{k-1}(x)$, we may write $g(x)=\hat{g}_{k}-Hp^{N}_{k-1}(x)$. Since
$p_{k-1}^{N}(x)=\sum_{i=0}^{k-1}-\frac{g(x)^{T}q_{i}}{q_{i}^{T}Hq_{i}}q_{i}.$
by Lemma 3.1, we have
$-\hat{g}_{k}^{T}g(x)=-\hat{g}_{k}^{T}(\hat{g}_{k}-Hp^{N}_{k-1}(x))=-\hat{g}_{k}^{T}\hat{g}_{k}-\frac{g(x)^{T}q_{k-1}}{q_{k-1}^{T}Hq_{k-1}}\hat{g}_{k}^{T}Hq_{k-1},$
where the orthogonality of $\hat{g}_{k}$ to $Hq_{i},i=0,\dots,k-2$ is used.
Then, $\beta_{k}$ and $\beta_{k-1}$ are given by
$\left(\!\begin{array}[]{ccccc}\hat{g}_{k}^{T}H\hat{g}_{k}&\hat{g}_{k}^{T}Hq_{k-1}\\ q_{k-1}^{T}H\hat{g}_{k}&q_{k-1}^{T}Hq_{k-1}\end{array}\!\right)\left(\!\begin{array}[]{c}\beta_{k}\\ \beta_{k-1}\end{array}\!\right)=\left(\!\begin{array}[]{c}-\hat{g}_{k}^{T}\hat{g}_{k}-\frac{g(x)^{T}q_{k-1}}{q_{k-1}^{T}Hq_{k-1}}\hat{g}_{k}^{T}Hq_{k-1}\\ -q_{k-1}^{T}g(x)\end{array}\!\right).$
Let $\beta_{k-1}=-\frac{g(x)^{T}q_{k-1}}{q_{k-1}^{T}Hq_{k-1}}+\gamma_{k-1}$, then this is reduced to
$\left(\!\begin{array}[]{ccccc}\hat{g}_{k}^{T}H\hat{g}_{k}&\hat{g}_{k}^{T}Hq_{k-1}\\ q_{k-1}^{T}H\hat{g}_{k}&q_{k-1}^{T}Hq_{k-1}\end{array}\!\right)\left(\!\begin{array}[]{c}\beta_{k}\\ \gamma_{k-1}\end{array}\!\right)=\left(\!\begin{array}[]{c}-\hat{g}_{k}^{T}\hat{g}_{k}\\ 0\end{array}\!\right),$
so that
$\left(\!\begin{array}[]{c}\beta_{k}\\ \gamma_{k-1}\end{array}\!\right)$ $=$ $-\frac{\hat{g}_{k}^{T}\hat{g}_{k}}{\hat{g}_{k}^{T}H\hat{g}_{k}q_{k-1}^{T}Hq_{k-1}-(\hat{g}_{k}^{T}Hq_{k-1})^{2}}\left(\!\begin{array}[]{ccccc}q_{k-1}^{T}Hq_{k-1}&-\hat{g}_{k}^{T}Hq_{k-1}\\ -q_{k-1}^{T}H\hat{g}_{k}&\hat{g}_{k}^{T}H\hat{g}_{k}\end{array}\!\right)\left(\!\begin{array}[]{c}1\\ 0\end{array}\!\right)$
$=$ $-\frac{\hat{g}_{k}^{T}\hat{g}_{k}}{\hat{g}_{k}^{T}H\hat{g}_{k}q_{k-1}^{T}Hq_{k-1}-(\hat{g}_{k}^{T}Hq_{k-1})^{2}}\left(\!\begin{array}[]{ccccc}q_{k-1}^{T}Hq_{k-1}\\ -q_{k-1}^{T}H\hat{g}_{k}\end{array}\!\right).$
The subspace Newton step to $\hat{x}_{k+1}$ is therefore formed by
$p^{N}_{k}(x)=\beta_{k}\hat{g}_{k}+\gamma_{k-1}q_{k-1}+\sum_{i=0}^{k-1}-\frac{g(x)^{T}q_{i}}{q_{i}^{T}Hq_{i}}q_{i}=\frac{1}{\hat{\sigma}_{k}}\left(-\hat{g}_{k}+\frac{\hat{g}_{k}^{T}Hq_{k-1}}{q_{k-1}^{T}Hq_{k-1}}q_{k-1}\right)+p_{k-1}^{N}(x)$
where
$\hat{\sigma}_{k}=-\frac{1}{\beta_{k}}=\frac{\hat{g}_{k}^{T}H\hat{g}_{k}q_{k-1}^{T}Hq_{k-1}-(\hat{g}_{k}^{T}Hq_{k-1})^{2}}{\hat{g}_{k}^{T}\hat{g}_{k}q_{k-1}^{T}Hq_{k-1}}.$
If we consider the case when $x=\hat{x}_{k}$, then $p^{N}_{k-1}(\hat{x}_{k})=0$. Therefore
$p^{N}_{k}(\hat{x}_{k})=\frac{1}{\hat{\sigma}_{k}}\left(-\hat{g}_{k}+\frac{\hat{g}_{k}^{T}Hq_{k-1}}{q_{k-1}^{T}Hq_{k-1}}q_{k-1}\right)=\hat{q}_{k},$
is parallel to the search direction obtained using CG, BFGS and memoryless BFGS; Newton scaled with respect to $\hat{x}_{k}$.
Given that information is known on how $H$ acts on the whole space $\mathcal{K}_{k+1}(g_{0},H)$, it is possible to calculate the subspace Newton step to $\hat{x}_{k+1}$ using the vectors $\hat{g}_{k},q_{k-1},\dots,q_{0}$. Building on this framework, we present a quasi-Newton subspace step under the assumption that we have incomplete information on the action of $H$ on the Krylov subspace $\mathcal{K}_{k+1}(g_{0},H)$.
4 Subspace quasi-Newton methods
To recover the subspace Newton step from $x\in x_{0}+\mathcal{K}_{k}(g_{0},H)$ to $\hat{x}_{k+1}$ as presented in the previous section, we require knowledge on how $H$ acts on the full subspace $K_{k+1}(g_{0},H)$. In particular sufficient information is needed to calculate the Newton scaling of the conjugate vectors $q_{0},\dots,q_{k-1}$ with respect to $x$ as well as the scalars $\beta_{k},\gamma_{k-1}$. Suppose instead we simply have a quasi-Newton approximation $B_{k}\succ 0$ such that $B_{k}q_{i}=Hq_{i}$ for $i=0,\dots,k-1$. Then the only term that is unknown is $\hat{g}_{k}^{T}H\hat{g}_{k}$, which must be replaced by the approximation $\hat{g}_{k}^{T}B_{k}\hat{g}_{k}$.
If such an approximation is used, it is possible to replace $H$ with $B_{k}$ and use the construction in Proposition 3.1 to generate a step $p_{k}(x)=q_{k}+p^{N}_{k-1}(x)$ where
$q_{k}=\frac{\hat{\sigma}_{k}}{\sigma_{k}}\hat{q}_{k}\text{for }\sigma_{k}=\frac{\hat{g}_{k}^{T}B_{k}\hat{g}_{k}q_{k-1}^{T}Hq_{k-1}-(\hat{g}_{k}^{T}Hq_{k-1})^{2}}{\hat{g}_{k}^{T}\hat{g}_{k}q_{k-1}^{T}Hq_{k-1}},$
taking into account that $B_{k}q_{i}=Hq_{i},i=0,\dots,k-1$. However, this construction requires explicit knowledge of the gradient $\hat{g}_{k}$. Additionally, $\sigma_{k}$ is undefined when $k=r$. To avoid this, we will build an approximation $B_{k}$ which acts on the Krylov subspace $\mathcal{K}_{k}(g_{0},H)$ for any $k\leq r$ in such a way that for any $x\in x_{0}+\mathcal{K}_{k}(g_{0},H)$, solving the system of equations $B_{k}p_{k}(x)=-g(x)$ yields a step $p_{k}(x)=q_{k}+p_{k-1}^{N}(x)$ where $q_{k}//\hat{q}_{k}$, without requiring knowledge of $\hat{g}_{k}$.
Lemma 4.1
For $k\leq r$, let $x\in x_{0}+\mathcal{K}_{k}(g_{0},H)$ and let $Q_{k-1}=(q_{k-1}\ q_{k-2}\ \dots\ q_{0})$ for $q_{i}//\hat{q}_{i},i=0,\dots,k-1$. Assume that $B_{k}$ is given by
$B_{k}=\sigma_{k}(I-Q_{k-1}(Q_{k-1}^{T}Q_{k-1})^{-1}Q_{k-1}^{T})+HQ_{k-1}(Q_{k-1}^{T}HQ_{k-1})^{-1}Q_{k-1}^{T}H,$ (4.1)
where $\sigma_{k}>0$. Then $B_{k}p_{k}=-g(x)$ has a unique solution given by
$p_{k}(x)$ $=q_{k}+p^{N}_{k-1}(x)$ (4.2)
where, for $k<r$,
$q_{k}=\frac{1}{\sigma_{k}}\left(-\hat{g}_{k}+\frac{\hat{g}_{k}^{T}\!Hq_{k-1}}{q_{k-1}^{T}Hq_{k-1}}q_{k-1}\right)=\frac{\hat{\sigma}_{k}}{\sigma_{k}}\hat{q}_{k}.$
for $\hat{\sigma}_{k}$ and $\hat{q}_{k}$ given by Proposition 3.1. As a consequence, when $\sigma_{k}=\hat{\sigma}_{k}$, the solution is the subspace Newton step $p^{N}_{k}(x)$ to $\hat{x}_{k+1}$. For $k=r$, $q_{r}=\hat{q}_{r}=0$ and the solution $p_{r}(x)=p^{N}_{r-1}(x)$ is independent of the choice of $\sigma_{r}$.
Proof. Since $\hat{g}_{k}$ is the gradient at $\hat{x}_{k}$, it must be orthogonal to the space spanned by the columns of $Q_{k-1}$. Additionally, by construction the term
$q_{k}=\frac{1}{\sigma_{k}}\left(-\hat{g}_{k}+\frac{\hat{g}_{k}^{T}\!Hq_{k-1}}{q_{k-1}^{T}Hq_{k-1}}q_{k-1}\right)$
is conjugate to the vectors $q_{0},\dots,q_{k-1}$ with respect to $H$ and belongs to $K_{k+1}(g_{0},H)$. Therefore $B_{k}q_{k}=-\hat{g}_{k}$.
By Lemma 3.1, $p^{N}_{k-1}(x)$ is a linear combination of the vectors $q_{0},\dots,q_{k-1}$. Therefore $B_{k}p^{N}_{k-1}(x)=Hp^{N}_{k-1}(x)$. Since $p^{N}_{k-1}(x)$ is the subspace Newton step to $\hat{x}_{k}$, the gradient at $x$ may be rewritten as $g(x)=\hat{g}_{k}-Hp^{N}_{k-1}(x).$
Applying $B_{k}$ to (4.2) therefore yields the result
$B_{k}p_{k}(x)=-\hat{g}_{k}+Hp^{N}_{k-1}(x)=-g(x).$
As $Q_{k-1}$ has full column rank, and $\sigma_{k}>0$, $B_{k}$ is strictly positive definite by Lemma A.2 and the solution $p_{k}(x)$ is therefore unique. For the case where $\sigma_{k}=\hat{\sigma}_{k}$, the solution is exactly $p^{N}_{k}(x)=\hat{q}_{k}+p^{N}_{k-1}(x)$.
When $k=r$, the gradient $\hat{g}_{r}=0$. Therefore, by definition, $q_{r}=\hat{q}_{r}=0$. For any $\sigma_{r}>0$,
$B_{r}p^{N}_{r-1}(x)=Hp^{N}_{r-1}(x)=-g(x),$
so $p^{N}_{r-1}(x)=p_{r}(x)$.
The approximation $B_{k}$ given by (4.1) requires a full basis of $\mathcal{K}_{k}(g_{0},H)$ to calculate the step $p_{k}(x)$. In the proof of Lemma 4.1 however, it is unnecessary for $B_{k}$ to act on each basis vector $q_{0},\dots,q_{k-1}$ as $H$ individually. The key components of the proof require only that $B_{k}$ acts as $H$ on $q_{k-1}$ and $p^{N}_{k-1}(x)$.
We now show that is suffices to replace the full basis of $\mathcal{K}_{k}(g_{0},H)$ by the vectors $q_{k-1}$ and $p^{N}_{k-1}(x)$ in the construction of a Hessian approximation $B_{k}(x)$. Then $B_{k}(x)$ need only act as $H$ on a subspace of dimension at most two, regardless of the size of $\mathcal{K}_{k}(g_{0},H)$.
Proposition 4.1
For $k\leq r$, let $x\in x_{0}+\mathcal{K}_{k}(g_{0},H)$, let $p^{N}_{k-1}(x)$ be the subspace Newton step to $\hat{x}_{k}$ and let $q_{k-1}//\hat{q}_{k-1}$. Define
$P_{k-1}=\begin{cases}\begin{bmatrix}p^{N}_{k-1}(x)&q_{k-1}\end{bmatrix},&\text{if }\mathop{\operator@font rank}\left(\begin{bmatrix}p^{N}_{k-1}(x)&q_{k-1}\end{bmatrix}\right)=2,\\[10.0pt] \begin{bmatrix}q_{k-1}\end{bmatrix},&\text{otherwise.}\end{cases}$
Assume that $B_{k}(x)$ is given by
$B_{k}(x)=\sigma_{k}(I-P_{k-1}(P_{k-1}^{T}P_{k-1})^{-1}P_{k-1}^{T})+HP_{k-1}(P_{k-1}^{T}HP_{k-1})^{-1}P_{k-1}^{T}H.$ (4.3)
Then $B_{k}(x)p_{k}(x)=-g(x)$ has a unique solution
$p_{k}(x)$ $=q_{k}+p^{N}_{k-1}(x)$ (4.4)
which coincides with the solution in Lemma 4.1 calculated by the approximation $B_{k}$ given by (4.1).
Proof. Since $k\leq r$, $q_{k-1}$ must be nonzero. Therefore $P_{k-1}$ has full column rank by definition, so $B_{k}(x)$ is well defined. Additionally, $B_{k}(x)$ is positive definite by Lemma A.2, so a solution to the equation $B_{k}(x)p_{k}(x)=-g_{k}(x)$ must exist and be unique.
Let
$q_{k}=\frac{1}{\sigma_{k}}\left(-\hat{g}_{k}+\frac{\hat{g}_{k}^{T}Hq_{k-1}}{q_{k-1}^{T}Hq_{k-1}}q_{k-1}\right).$
By construction, $q_{k}$ is conjugate to $\mathcal{K}_{k}(g_{0},H)$ with respect to $H$. By Lemma 3.1 this implies that $q_{k}^{T}Hp^{N}_{k-1}(x)=0$. Therefore $B_{k}q_{k}=-\hat{g}_{k}$. Additionally, by design $B_{k}(x)p^{N}_{k-1}(x)=Hp^{N}_{k-1}(x)$. Applying $B_{k}$ to (4.4) yields
$B_{k}(x)p_{k}(x)=-\hat{g}_{k}+Hp^{N}_{k-1}(x)=-g(x)$
which follows from $p^{N}_{k-1}(x)$ being the subspace Newton step to $\hat{x}_{k}$.
Therefore $p_{k}(x)=q_{k}+p^{N}_{k-1}(x)$ is the unique solution to the system of equations  $B_{k}(x)p_{k}(x)=-g_{k}(x)$ and coincides with the solution calculated by the approximation $B_{k}$ given by (4.1).
For any point $x$ in the affine space $x_{0}+\mathcal{K}_{k}(g_{0},H)$ we can now construct a Hessian approximation using the Newton subspace step $p^{N}_{k-1}(x)$ and a vector $q_{k-1}//\hat{q}_{k-1}$. This Hessian approximation along with the gradient yields a quasi-Newton step that is a combination of $p_{k-1}^{N}(x)$ and $q_{k}//\hat{q}_{k}$. We have shown that there is a unique value $\hat{\sigma}_{k}$ such that the resulting step is exactly the subspace Newton step $p^{N}_{k}(x)$ to $\hat{x}_{k+1}$. However, even if the resulting step is not the Newton subspace step, the point $x+\alpha p(x)$ for any step size $\alpha$ belongs to the affine space $x_{0}+\mathcal{K}_{k+1}(g_{0},H)$. We may use this property to create an iterative method which successively generates quasi-Newton steps regardless of step size.
5 An exact quasi-Newton method without exact line search
If we are given an arbitrary point $x\in x_{0}+\mathcal{K}_{k-1}(g_{0},H)$, then generating the subspace quasi-Newton step as described in Proposition 4.1 requires explicit knowledge of the direction of the subspace Newton step $p^{N}_{k-1}(x)$ from $x$ to $\hat{x}_{k}$ as well as the direction $\hat{q}_{k-1}$. In addition it is also vital that we have information on how $H$ acts on these vectors to build the Hessian approximation $B_{k}$ given by (4.3). We propose an iterative learning algorithm which generates subspace quasi-Newton steps regardless of choice of step sizes. After a finite number of iterations, the algorithm will generate the Newton step, which will lead to termination if a unit step is used. The step size $\alpha_{k}$ can be chosen freely. This method is described in Algorithm 1.
Theoretically, if we have explicit knowledge of $H$ then even a zero step will lead to generation of the Newton step as iteratively new conjugate directions are calculated and then Newton scaled with respect to the fixed point. Without explicit knowledge of $H$ it is vital to take a nonzero step in order to evaluate a new gradient. It is then possible to use a difference of gradients to learn how $H$ acts on the step taken. This is described in Lemma A.1.
Algorithm 1 An exact quasi-Newton method for solving (QP)
$x_{0}\in\mathbb{R}^{n},p^{N}_{-1}=0,\sigma_{0}>0$
$k\leftarrow 0$;
$g_{k}\leftarrow Hx_{k}+c$;
while $\|g_{k}\|\neq 0$ do
     if $k=0$ then
         $B_{k}=\sigma_{k}I$;
     else
         Choose some $\sigma_{k}>0$;
         $B_{k}=\sigma_{k}(I-P_{k-1}(P_{k-1}^{T}P_{k-1})^{-1}P_{k-1}^{T})+HP_{k-1}(P_{k-1}^{T}HP_{k-1})^{-1}P_{k-1}^{T}H$;
     end if
     Solve $B_{k}p_{k}=-g_{k}$;
     Choose step size $\alpha_{k}$;
     $x_{k+1}\leftarrow x_{k}+\alpha_{k}p_{k}$;
     $g_{k+1}\leftarrow g_{k}+\alpha_{k}Hp_{k}$;
     if $\|g_{k+1}\|=0$ then
         break
     end if
     $q_{k}\leftarrow p_{k}-p^{N}_{k-1}$;
     if $\|q_{k}\|\neq 0$ then
         $p^{N}_{k}\leftarrow\left(-\frac{g_{k}^{T}q_{k}}{q_{k}^{T}Hq_{k}}-1\right)q_{k}+(1-\alpha_{k})p_{k}$;
[…]
              $P_{k}=\begin{bmatrix}p^{N}_{k}&q_{k}\end{bmatrix}$;
         else
              $P_{k}=\begin{bmatrix}q_{k}\end{bmatrix}$;
         end if
     else
         $p^{N}_{k}=(1-\alpha_{k})p^{N}_{k-1}$;
         $P_{k}=\begin{bmatrix}p^{N}_{k}\end{bmatrix}$;
     end if
     $k\leftarrow k+1$;
end while
Theorem 5.1
Let $x_{0}$ be a fixed initial point and let $x_{i},i=0,\dots k$ be generated by Algorithm 1 where $\alpha_{i}$ are arbitrarily chosen for $i<k$ and $\sigma_{i}>0$ for $i\leq k$. When $k\leq r$, the search direction $p_{k}$ generated by the algorithm is the search direction $p_{k}(x_{k})$ defined in Proposition 4.1 as given by (4.4). Additionally, for $k\geq r$, the search direction $p_{k}$ generated by Algorithm 1 is precisely the Newton step so the algorithm will terminate if $\alpha_{k}=1$.
Proof. When $k<r$, to show that the search direction $p_{k+1}$ generated by Algorithm 1 is exactly $p_{k+1}(x_{k+1})$ given by (4.4), it is sufficient to show that $q_{k}//\hat{q}_{k}$ and that $p_{k}^{N}=p_{k}^{N}(x_{k+1})$. We proceed by induction. For $k=0$, clearly $\sigma_{0}q_{0}=-g(x_{0})$ and $p_{0}^{N}$ is Newton scaled with respect to $x_{1}$.
Assume that $q_{k-1}//\hat{q}_{k-1}$ and $p^{N}_{k-1}=p^{N}_{k-1}(x_{k})$, then by Proposition 4.1 $p_{k}$ is exactly $p_{k}(x_{k})$ which of the form $p_{k}=q_{k}+p^{N}_{k-1}$, therefore $q_{k}//\hat{q}_{k}$.
As the search directions generated by the algorithm coincides with those given in Proposition 4.1 for $i\leq k$, it is straightforward to show that $x_{k+1}\in x_{0}+\mathcal{K}_{k+1}(g_{0},H)$. Using the formulation for the subspace Newton step in Lemma 3.1, we have
$p_{k}^{N}(x_{k+1})$ $=-\sum_{i=0}^{k}\frac{g(x_{k+1})^{T}q_{i}}{q_{i}^{T}Hq_{i}}q_{i}$
$=-\sum_{i=0}^{k}\frac{(g_{k}+\alpha_{k}Hq_{k}+\alpha_{k}Hp_{k-1}^{N}(x_{k}))^{T}q_{i}}{q_{i}^{T}Hq_{i}}q_{i}$
$=\left(-\frac{g_{k}^{T}\!q_{k}}{q_{k}^{T}\!Hq_{k}}-\alpha_{k}\right)q_{k}-\sum_{i=0}^{k-1}\frac{g(x_{k}+\alpha_{k}p_{k-1}^{N}(x_{k}))^{T}q_{i}}{q_{i}^{T}Hq_{i}}q_{i}$
$=\left(-\frac{g_{k}^{T}\!q_{k}}{q_{k}^{T}\!Hq_{k}}-\alpha_{k}\right)q_{k}+p_{k-1}^{N}(x_{k}+\alpha_{k}p_{k-1}^{N}(x_{k}))$
$=\left(-\frac{g_{k}^{T}\!q_{k}}{q_{k}^{T}\!Hq_{k}}-\alpha_{k}\right)q_{k}+(1-\alpha_{k})p_{k-1}^{N}(x_{k})$
$=\left(-\frac{g_{k}^{T}\!q_{k}}{q_{k}^{T}\!Hq_{k}}-1\right)q_{k}+(1-\alpha_{k})p_{k},$ (5.1)
taking the conjugacy of $q_{k}$ with $p^{N}_{k-1}(x_{k}),q_{k-1},\dots,q_{0}$ into account, as well as the property of the subspace Newton step
$p^{N}_{k-1}(x_{k}+\alpha_{k}p^{N}_{k-1}(x_{k}))=\hat{x}_{k}-x_{k}-\alpha_{k}p^{N}_{k-1}(x_{k})=(1-\alpha_{k})p^{N}_{k-1}(x_{k}).$
Since $q_{k}$ is nonzero, $B_{k+1}$ is constructed by the algorithm exactly as $B_{k+1}(x_{k+1})$ is in Proposition 4.1. Therefore $p_{k+1}$ is the search direction $p_{k+1}(x_{k+1})$ given by (4.4).
As a consequence of Proposition 4.1, $p_{r}=p^{N}_{r-1}(x_{r})$ is the Newton step and $q_{r}=\hat{q}_{r}=0$. Therefore for any step $\alpha_{r}$, $p^{N}_{r}=(1-\alpha_{r})p^{N}_{r-1}$ is the Newton step from $x_{r+1}$. Proceeding again using induction, for any $k>r$, if $p^{N}_{k-1}$ is the Newton step and $q_{k}=0$, then $B_{k}$ defined by the algorithm will generate the Newton step. The algorithm will therefore terminate if $\alpha_{k}=1$.
As stated in Theorem 5.1, the search direction $p_{k}$ generated by Algorithm 1 for $k\geq n$ is precisely the Newton step. Therefore if a unit step size is chosen, the algorithm will terminate. It is however possible to terminate in exactly $r$ iterations if and only if the Hessian approximation is built using the unique value $\hat{\sigma}_{r-1}$ stated in Proposition 3.1.
Corollary 5.1
For the unique value $\hat{\sigma}_{r-1}$ stated in Proposition 3.1, Algorithm 1 will generate the Newton step $p_{r-1}=p^{N}_{r-1}(x_{r-1})$, and the algorithm will terminate at iteration $r$ if $\alpha_{r-1}=1$. For all other value of $\sigma_{r-1}$ the algorithm will terminate only when $\alpha_{k}=1$ for some $k\geq r$.
Proof. As stated in Theorem 5.1, for $k\leq r$ the search direction $p_{k}$ coincides with the search direction $p_{k}(x_{k})$ from Proposition 4.1. Therefore, for the unique value of $\hat{\sigma}_{r-1}$ stated in Proposition 3.1
$p_{r-1}=\frac{\hat{\sigma}_{r-1}}{\hat{\sigma}_{r-1}}\hat{q}_{r-1}+p^{N}_{r-2}(x_{r-1})=p^{N}_{r-1}(x_{r-1}).$
Hence the algorithm will terminate at iteration $r$ if $\alpha_{r-1}=1$. For all other values of $\sigma_{r-1}$, the search direction $p_{r-1}$ cannot be parallel to $p^{N}_{r-1}(x_{r-1})$ as $\hat{q}_{r-1}\neq 0$. Therefore the algorithm cannot terminate. Applying Theorem 5.1, for $k\geq r$ the search direction $p_{k}=p^{N}_{r-1}(x_{k})$ is the Newton step, so the algorithm will terminate only if $\alpha_{k}=1$.
[…]
This method will generate the Newton step in a finite number of iterations regardless of step sizes chosen. The action of $B_{k}$ on the correct subspace of two vectors is the vital component which yields the subspace quasi-Newton step of Proposition 4.1. As shown in Lemma A.3 solving $B_{k}p_{k}=-g_{k}$ can be done by solving two systems of linear equations where the matrix is $P_{k}^{T}HP_{k}$, whose dimension is at most $2\times 2$.
6 Conclusion
In this work we explore exactness in quasi-Newton methods for minimizing a strictly convex quadratic function without the need for the use of exact line search. With knowledge of how the Hessian acts on at most two vectors in the space $\mathcal{K}_{k}(g_{0},H)$, it is possible to generate a quasi-Newton step which is composed of a subspace Newton step to $\hat{x}_{k}$ and a vector parallel to the difference of minimizers $\hat{q}_{k}=\hat{x}_{k+1}-\hat{x}_{k}.$ Using this framework, we have presented an iterative learning algorithm which successively generates such quasi-Newton steps for expanding affine spaces, regardless of step size used.
The algorithm which has been presented generates parallel search directions to that of CG, BFGS and memoryless BFGS when exact line search is used. In the case when exact line search is used and $\sigma_{k}=1$ always, the Hessian approximation $B_{k}$ generated by Algorithm 1 is in fact precisely the approximation that is built using memoryless BFGS. Algorithm 1 can therefore be viewed as a generalization of memoryless BFGS.
The results presented in this work are meant to deepen the theoretical understanding of the action of quasi-Newton methods on quadratic problems. The algorithm presented is a conceptual algorithm to explore the conditions under which it is possible to achieve finite termination on (QP) assuming exact arithmetic. We also hope that our work can lead further research to get a deeper understanding of the behavior of quasi-Newton methods on general nonlinear problems.
Appendix A Appendices
A.1 A first order formulation of Algorithm 1
The construction of $B_{k+1}$ in Algorithm 1 requires information on how $H$ acts on the vectors $q_{k}$ and $p^{N}_{k}$. When a nonzero step is taken at each iteration, this information can be gathered using a difference of gradients. Explicit knowledge of the Hessian $H$ is unnecessary.
Lemma A.1
For $k<r$, if $\alpha_{k}\neq 0$ is chosen in Algorithm 1, then the action of $H$ on $q_{k}$ can be calculated by
$Hq_{k}=\frac{1}{\alpha_{k}}(g_{k+1}-g_{k})-Hp^{N}_{k-1}$ (A.1)
where $Hp^{N}_{k-1}=B_{k}p^{N}_{k-1}.$ Using this formulation, $p^{N}_{k}$ and $Hp^{N}_{k}$ can be recovered by
$p^{N}_{k}=\left(-\frac{\alpha_{k}g_{k}^{T}q_{k}}{q_{k}^{T}(g_{k+1}-g_{k})}-1\right)q_{k}+(1-\alpha_{k})p_{k}$ (A.2)
$Hp^{N}_{k}=\left(-\frac{\alpha_{k}g_{k}^{T}q_{k}}{q_{k}^{T}(g_{k+1}-g_{k})}-1\right)Hq_{k}+\left(\frac{1}{\alpha_{k}}-1\right)(g_{k+1}-g_{k}).$ (A.3)
When $k\geq r$, $p^{N}_{k}=(1-\alpha_{k})p_{k}$, so
$Hp^{N}_{k}=\left(\frac{1}{\alpha_{k}}-1\right)(g_{k+1}-g_{k}).$ (A.4)
Proof. The step generated by Algorithm 1 is of the form $p_{k}=q_{k}+p^{N}_{k-1}$. Assuming a nonzero step size $\alpha_{k}$ is used so that $x_{k+1}=x_{k}+\alpha_{k}p_{k}$, evaluating the gradient $g_{k+1}$ yields
$\frac{1}{\alpha_{k}}(g_{k+1}-g_{k})=Hp_{k}=Hq_{k}+Hp^{N}_{k-1}.$
Given that the matrix $B_{k}$ acts as $H$ on $p^{N}_{k-1}$, the formulation of $Hq_{k}$ in (A.1) follows directly.
As $p^{N}_{k-1}\in\mathcal{K}_{k}(g_{0},H)$, it is conjugate to $q_{k}$ with respect to $H$. Therefore we may replace the term $q_{k}^{T}Hq_{k}$ by
$q_{k}^{T}Hq_{k}$ $=q_{k}^{T}\left(\frac{1}{\alpha_{k}}(g_{k+1}-g_{k})-Hp^{N}_{k-1}\right)$
$=\frac{1}{\alpha_{k}}q_{k}^{T}(g_{k+1}-g_{k})$
in (5.1) to obtain the expression (A.2). The formulations for $Hp^{N}_{k}$ in (A.3) and (A.4) follow immediately.
Using the expressions for $p^{N}_{k},Hq_{k}$ and $Hp^{N}_{k}$ given by the lemma above means that Algorithm 1 can be implemented using first order information only, with the same termination guarantees.
A.2 Some linear algebra results
In this appendix, a couple of linear algebra results are given. The first lemma shows that the quasi-Newton matrices generated are positive definite.
Lemma A.2
Let
$B=\sigma(I-P(P^{T}P)^{-1}P^{T})+HP(P^{T}HP)^{-1}P^{T}\!H,$
where $H$ is a symmetric positive definite $n\times n$ matrix, $P$ is an $n\times k$ matrix of rank $k$, and $\sigma$ is a positive constant. Then, $B$ is positive definite.
Proof. We have $B=B_{1}+B_{2}$, for
$B_{1}=\sigma(I-P(P^{T}P)^{-1}P^{T})\text{and}B_{2}=HP(P^{T}HP)^{-1}P^{T}\!H.$
[…]
The following lemma shows a way to solve with the low rank quasi-Newton matrix.
Lemma A.3
Let
$B=\sigma(I-P(P^{T}P)^{-1}P^{T})+HP(P^{T}HP)^{-1}P^{T}\!H,$
where $H$ is a symmetric positive definite $n\times n$ matrix, $P$ is an $n\times k$ matrix of rank $k$, and $\sigma$ is a positive constant. Then, $Bp=-g$ if and only if
$P^{T}\!HP\beta$ $=$ $-P^{T}\!g,$
$P^{T}\!HP\delta$ $=$ $-P^{T}\!Hg-(\sigma P^{T}\!HP+P^{T}\!H^{2}P)\beta,$
$p$ $=$ $-\frac{1}{\sigma}\left(g+P\delta+HP\beta\right).$
Proof. Lemma A.2 shows that $B$ is positive definite, so that $p$ is well defined and unique. In addition, we have
$(\sigma(I-P(P^{T}P)^{-1}P^{T})+HP(P^{T}HP)^{-1}P^{T}\!H)p=-g$
if and only if
$\left(\!\begin{array}[]{ccc}\sigma I&P&HP\\ P^{T}&\frac{1}{\sigma}P^{T}\!P&0\\ P^{T}\!H&0&-P^{T}\!HP\end{array}\!\right)\left(\!\begin{array}[]{c}p\\ \delta\\ \beta\end{array}\!\right)=-\left(\!\begin{array}[]{c}g\\ 0\\ 0\end{array}\!\right)$
for some $k$-dimensional vectors $\delta$ and $\beta$. Elimination of $p$ in the second and third block of equations gives
$\left(\!\begin{array}[]{ccc}\sigma I&P&HP\\ 0&0&-\frac{1}{\sigma}P^{T}\!HP\\ 0&-\frac{1}{\sigma}P^{T}\!HP&-P^{T}\!HP-\frac{1}{\sigma}P^{T}\!H^{2}P\end{array}\!\right)\left(\!\begin{array}[]{c}p\\ \delta\\ \beta\end{array}\!\right)=\left(\!\begin{array}[]{c}-g\\ \frac{1}{\sigma}P^{T}g\\ \frac{1}{\sigma}P^{T}\!Hg\end{array}\!\right).$
We may therefore solve in turn
$P^{T}\!HP\beta$ $=$ $-P^{T}\!g,$
$P^{T}\!HP\delta$ $=$ $-P^{T}\!Hg-(\sigma P^{T}\!HP+P^{T}\!H^{2}P)\beta,$
$p$ $=$ $-\frac{1}{\sigma}\left(g+P\delta+HP\beta\right),$
completing the proof.
This means solving with $B$ can be done by solving two systems of dimension equal to the number of columns in $P$, in our low-rank setting two by two.
[…]